Pb(NO3)2 + 2 NaI à PbI2 + 2 NaNO

If 100 grams of lead (II) nitrate were used in the reaction, how many grams of sodium iodide were used for a complete reaction? 

1
Expert's answer
2022-03-08T11:03:04-0500

Pb(NO3)2 + 2 NaI = PbI2 + 2 NaNO3

According to the equation, n (NaI) = 2 x n (Pb(NO3)2).

n = m / M

M (Pb(NO3)2) = 331.2 g/mol

M (NaI) = 149.9 g/mol

n (Pb(NO3)2) = m / M = 100 / 331.2 = 0.3 mol

n (NaI) = 2 x n (Pb(NO3)2) = 2 x 0.3 = 0.6 mol

m (NaI) = n x M = 0.6 x 149.9 = 90 g